OpenStudy (anonymous):

Need help in the following question: The area of a rectangle is 30 square feet. The length of the rectangle is 4 feet shorter than double the width. Find the length and the width.

OpenStudy (anonymous):

Area of rectangle = (length) x (width) = ( (2 x width) - 4) x (width) solve for width then plug back in to solve for length

OpenStudy (anonymous):

im not understanding, sorry.

OpenStudy (calculusxy):

Lets put this into an equation, so that we can find the length first: w^2-4=length

OpenStudy (anonymous):

okay sounds good !

OpenStudy (calculusxy):

Now we can use guess and check to replace the 'w' variable.

OpenStudy (anonymous):

plug in any number that will equal the are of the rectangle -30 square feet?

OpenStudy (anonymous):

*area

OpenStudy (calculusxy):

Actually that will equal to the length. Becuase after the length, we can easily find the width

OpenStudy (anonymous):

oh so how can i do that ?

OpenStudy (calculusxy):

Wait, I am trying to figure it out.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

okay thank you i'll check this and figure what info. i have and don't have

OpenStudy (calculusxy):

No problem :)

OpenStudy (anonymous):

Use what you learned from calculusxy that the length is = 2 times the width - 4 so in equation form it would look like this with "w" representing the "width"\[length = (2w-4)\] my earlier equation represented the definition of the area of rectangle\[Area = (length)(width)\] plug in your length which was represented up there to get this equation\[30 = (2w-4)(w)\] simplify it to this\[30 = 2w^2-4w\] then to this\[30 = 2(w^2-2w)\]complete the square for w^2-2w to get:\[30+2 = 2(w^2-2w+1)\]simplify it to this\[32 = 2(w-1)^2\]solving for "w"\[16 = (w-1)^2\]\[4 = w-1\]\[w =3\] now plug that into your original equation\[30 = (length)(3)\] then solve for length\[length = 10\]